FEL Upgrade Project Weekly Brief
November 26-30, 2001


Highlights:
On Nov. 19th the IR Demo FEL delivered its last photons prior to being decommissioned for start of the
installation process for the IR Demo Upgrade.  For its last weekend of operation the IR Demo was used for
four concluding runs for laser production of carbon nanotubes.  The nanotube team reports that they
approximately doubled their inventory of nanotube material.

Immediately after the final run, the decommissioning and dis-assembly process began.  All components of the
Demo were barcoded for inventory and specific components which are going to be re-used in the Upgrade were
specifically marked.  By the end of this week, all cabling in the 4F and 5F regions of the machine had been
disconnecting and the backleg vacuum system had been removed.  The shielding blocks were removed from the
access door to the accelerator vault for removal and installation
of major hardware.

As a second shift activity , the SRF team performed high power pulsed processing on the second cavity in the injector cryounit improving its maximum gradient by at least 10%. We thank Charlie Reece for organizing and overseeing this activity.

Management:
Following the November semiannual project review, we are taking the committee's advice and carefully
analyzing all budget categories for the remaining tasks in the Upgrade project to identify options for cost savings
or scope reduction.  These steps are also necessary since it appears highly likely that the FY02 DOD budget
allocations (that could provide some relief to our tight budget situation) will delayed well into the fiscal year.

The subcontract for participation by the Aerospace Corporation in the UV FEL project was forwarded to the
DOE Site Office for approval.

This week we had a visiting team from AES, Inc. for discussions of the joint work statement for the start-up of
the 100 mA injector project.  This start-up is possible because AES was recently awarded a Phase 2 SBIR from
SMDC (with supporting funds from NAVSEA) to design and  fabricate key components of a 750 MHz injector
cryounit compatible with the JLab FEL photogun.

We also had visitors from the Indian Head NSWC (William Walker and Walter Ott) who are interested in
potential collaboration with JLab and SURA universities, and from LBNL to discuss potential collaboration on
next generation x-ray sources. Finally, we had visitors from Dahlgren Labs and Northrop Grumman to discuss
cryoplant designs for high average power FELs.  A busy week for visitors.

WBS 4 (Injector):
Baked the FET system with an implanted electrode installed, testing will start next week. The parts for the gun
chamber have been completed by the shop, they will be cleaned prior to welding.

WBS 6 (RF):
Zone 2 - The 5 watt RF Amplifiers were received, but are not yet installed and tested.

Zone 3 - No change, essentially finished.

Zone 4 - No change, essentially finished.  Position 8 klystron, circulator, etc. will be tested as a system this
afternoon.

Quarter - High Peak Power Processing (HPPP) was run on cavity 3.  The field emission of the cavity was
improved.  Additional processing will be performed today and several days next week at higher power levels.

100 kW Klystron - The LCW plumbing of the klystron cart is nearly finished.  The flow rates will be checked
with the existing LCW taps to determine the power levels for preliminary testing of the klystrons.

Quarter HVPS - The circuit design of the HVPS is nearly complete with the purchasing of the HV capacitors
and resistors next week.  The schematic will be updated and an informal review held.  In order to remove the
existing transformer, inductor, and variac, the HVPS's will have to be removed from the FEL.  The FEL gallery
floor will not safely support a fork-lift truck.

WBS 8 (Instrumentation):
Cabling has been removed from the 4F and 5F regions in support for removing the components in these areas.

All components in this area have also been barcoded for tracking purposes.

The individual labs are being re-organized and re-stocked for storage purposes until they will be used again.

A new VME interface circuit board design is nearing completion.

WBS 9 (Transport):
Dipoles
Optical Chicane Dipoles (GW)
o  We inspected the initial winding of the first coil as well as bend up fixtures and the potting fixture
    machining at Magnet Enterprises International last week.  Potting will be observed next week by an
    epoxy expert, Jeffrey Hubrig of Innovation Services.
Injector Dipoles (DU/DV)
o  DULY is almost finished revising small portions of the final assemblies.  The specifications are
    being written to match the drawings and revised with what we have learned from our procurements
    of magnets to date.
Arc Dipoles (GY, GX, GQ)
o  The magnetic model of the Bend and Reverse Bend Dipoles (GX & GQ) is being exercised to is
    being exercised to increase the width of the of the good field region per David Douglasí
    specification. The GX is nearly through and the GQ is starting.  The drawings will be revised when
    the adjustment is over.
o  The bids for the 180° Dipoles (DY) are due on Monday.
Magnet Measurement
o  The parts for the dipole measurement stand are signed off
o  We continued work on the probe that will measure the sextupole and the octupole.
Quadrupoles
3 inch quad (QX)
o  The core vender, New England Techni-Coil is making the final adjustment to the 45 degree chamfer
    on the pole edge of the first article core.
o  We resolved the non-problem of turn to turn shorts in the coils by unwrapping the copper tubing of
    the cooling circuits.  The wrapped configuration was giving false indications of shorts.
o  A second potential problem is being resolved by a thermal test.  Sporadic brazing of the tubes to
    cooling plates on the initial series of coils is being compared in a thermal test with coils that have
    continuous brazing.
o  A shipment of 64 fully qualified coils is being readied at Magnet Enterprises International.
Trim Quad (QT)
o  The bids were received and evaluated.  The low bidder is being asked to detail a cost breakdown
    because of the spread in the quotations.
Sextupole (SF)
o  DULY Research continues to model the sextupole.
Octupole (OT)
o  Work on this magnet is on furlough.
Corrector Dipoles (DB, DJ)
o  Milhous Co. is about 6 shy of delivering the entire order DB and DJ coils.
Beam Line and Vacuum
o  Dave Waldman chaired the kick-off meeting to start the task order with AES to design the Arc
    Chambers.
o  Several sample chambers with the thicker plates necessary for the upgrade have been made to test
    out the magnetic properties.
o  Design work on girders and stands continued. All details of the region between the cryomodules are
    now complete.
o  All return leg girders have been received.
o  All the new three leg pedestals are now out for bid.

WBS 11 (Optics):
Progress for the Upgrade occurred in the following areas:

With assistance from K. Capek (Accel. Div.), an ultraviewer was assembled and is undergoing tests in User
Lab 6. We have "mocked up" one-half of the optical cavity beam path in User Lab 2 and confirmed calculations
of beam sizes at various points in the path. This is an important check, as we are close to releasing drawings for
the OCMMS (optical cavity mirror metrology system) beamline components. Before the holidays the OCMMS
drawings were reviewed, marked up, and returned for revision. We are scheduling an internal review for early
next week.

We have received the following parts for the optical cavities: ball bearings for the leadscrew, flex pivots for the
yaw axes, and flex pivots for the pitch axes. The bearings are being cleaned for UHV service.

Detailing concentrated on components from the "big C" (the yaw axis yoke) through the mirror holders. We
have identified a vendor for sputter-coating the leadscrews and are going out for procurement.

Meetings continued with sales representatives of ultrafast lasers. This laser will be used as a FEL cavity
dynamical diagnostic, and for diagnostics of the e-beam, using EO sampling. Further meetings were held on
procurement of a new drive laser.

Operations/Decommissioning activities:
On Friday, 16 Nov, mirrors were switched from 1.6 micron operation to 3 micron operation. This allowed the
nanotube collaboration to try some new targets. We provide support, in the way of a vacuum pump, to Siemens
Automotive. We also supported the SNS project by providing optical inspection metrology for their cavity
flanges. We documented, removed, and disassembled the THz cross-correlation experiment. Optical transport
mirror cans in the accelerator vault, as well as one OBPM, were removed. All orientations were documented for later metrology. Hardware loaned to users in User Lab 1, 3, & 6 and the X-ray experiment was collected, disassembled, and stored.